OpenStudy (anonymous):

change the base from 5 to e: log base 5 (x) then find the derivative

OpenStudy (unklerhaukus):

\[\large\log_bx=\frac{\log_ex}{\log_eb}=\frac{\ln x}{\ln b}\]

OpenStudy (anonymous):

whats the derivative of lne

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

lne? \(\ln e=\log_ee=1\) the derivative of a constant is zero

OpenStudy (ankitshaw):